OpenStudy (anonymous):

can someone help me out please with these questions please:1;find a general lineal equation (Ax+By+C=0) of the straight line that has the indicated properties,and sketck each line please. 1;PASSes through(-5,5) and has slope of -1/2. 2:sasses through(-3,-4)and(-2,-8). 3:is a vertical and passes through(-1,-1).

OpenStudy (anonymous):

1) you have the slope with a point on the line, thus the equation of the line is \[\frac{ y-5 }{ x-(-5) }=\frac{ -1 }{ 2 }\]\[=2y-10=-x-5\]\[=2y+x-5=0\]

OpenStudy (anonymous):

2) you have two points that are on the line, thus \[\frac{ y-(-4) }{ x-(-3) }=\frac{ -8-(-4) }{ -2-(-3) }\]then simplify it.

OpenStudy (anonymous):

3) For any vertical line that pass through a point (a, b), the equation of the line will be\[x=a\]thus the answer here is \[\Huge x=-1\]